# QSpec: Speculative Decoding with Complementary Quantization Schemes

### v1

@@ 4.3 Acceleration Evaluation
3B11footnotemark: 1 W4A16 8 326.1 360.2 429.0 392.4 434.7 401.7
16 529.1 653.8 820.3 673.2 882.2 772.9
32 679.2 906.8 1261.4 848.4 1439.3 1103.6
QSpec 8 501.5 (1.54$\times$) 537.2 (1.49$\times$) 640.4 (1.49$\times$) 593.1 (1.51$\times$) 646.6. (1.49$\times$) 592.7 (1.48$\times$) 1.50$\times$
16 731.9 (1.38$\times$) 837.8 (1.28$\times$) 863.2 (1.17$\times$) 875.9 (1.30$\times$) 1081.4 (1.23$\times$) 945.0 (1.22$\times$) 1.24$\times$
32 950.3 (1.40$\times$) 1175.8 (1.30$\times$) 1371.0 (1.09$\times$) 1052.3 (1.24$\times$) 1645.9 (1.14$\times$) 1347.6 (1.22$\times$) 1.23$\times$
Avg. 1.44$\times$ 1.36$\times$ 1.25$\times$ 1.35$\times$ 1.29$\times$ 1.31$\times$ 1.33$\times$
7B W4A16 8 126.8 144.1 165.0 170.4 177.1 157.0
16 213.1 267.2 314.9 344.6 358.6 300.8
32 257.6 347.1 409.7 478.9 509.1 402.0
QSpec 8 203.7 (1.61$\times$) 239.9 (1.66$\times$) 234.8 (1.42$\times$) 281.5 (1.65$\times$) 274.4 (1.55$\times$) 241.4 (1.54$\times$) 1.57$\times$
16 312.3 (1.47$\times$) 377.6 (1.41$\times$) 380.1 (1.21$\times$) 459.6 (1.33$\times$) 455.2 (1.27$\times$) 379.5 (1.26$\times$) 1.33$\times$
32 496.2 (1.54$\times$) 488.5 (1.41$\times$) 473.4 (1.16$\times$) 620.2 (1.30$\times$) 633.1 (1.24$\times$) 498.3 (1.24$\times$) 1.31$\times$
Avg. 1.54$\times$ 1.50$\times$ 1.26$\times$ 1.43$\times$ 1.35$\times$ 1.35$\times$ 1.40$\times$
8B W4A16 8 121.8 131.2 155.2 153.4 163.8 152.4
16 210.3 247.0 300.7 293.5 365.6 311.2
32 277.1 355.3 425.1 398.7 619.1 486.5
QSpec 8 191.7 (1.57$\times$) 200.4 (1.53$\times$) 214.4 (1.38$\times$) 230.1 (1.50$\times$) 241.0 (1.47$\times$) 220.6 (1.45$\times$) 1.48$\times$
16 294.2 (1.40$\times$) 333.4 (1.35$\times$) 334.1 (1.11$\times$) 373.0 (1.27$\times$) 431.7 (1.18$\times$) 373.4 (1.20$\times$) 1.25$\times$
32 368.8 (1.33$\times$) 447.5 (1.26$\times$) 478.1 (1.12$\times$) 484.2 (1.21$\times$) 687.3 (1.11$\times$) 564.1 (1.16$\times$) 1.20$\times$
Avg. 1.43$\times$ 1.38$\times$ 1.21$\times$ 1.33$\times$ 1.25$\times$ 1.27$\times$ 1.31$\times$
13B11footnotemark: 1 W4A16 8 74.0 85.1 103.7 100.5 104.1 92.3
16 128.6 163.0 185.8 177.7 222.8 173.1
32 195.1 206.9 323.7 327.8 330.1 241.6
QSpec 8 127.8 (1.73$\times$) 148.6 (1.75$\times$) 173.4 (1.67$\times$) 180.8 (1.80$\times$) 173.5 (1.68$\times$) 150.2 (1.63$\times$) 1.71$\times$
16 194.7 (1.51$\times$) 235.4 (1.44$\times$) 285.9 (1.54$\times$) 292.5 (1.65$\times$) 288.7 (1.30$\times$) 222.9 (1.29$\times$) 1.45$\times$
32 247.1 (1.27$\times$) 307.4 (1.49$\times$) 399.9 (1.24$\times$) 435.3 (1.33$\times$) 407.1 (1.23$\times$) 323.3 (1.34$\times$) 1.31$\times$
Avg. 1.50$\times$ 1.56$\times$ 1.48$\times$ 1.59$\times$ 1.40$\times$ 1.42$\times$ 1.49$\times$
QSpec exhibits a substantial efficiency boost compared to W4A16. In Table 4, we present the token generation throughput for both QSpec and W4A16 across different model sizes, quantization configurations, and batch sizes on diverse datasets. On average, QSpec achieves a throughput increase of 1.38×1.38$\times$1.38 × over W4A16 across all settings, with a peak improvement of 1.80×1.80$\times$1.80 ×.
Speedup does not show an evident correlation with task difficulty. As shown at the bottom of Table 4, we calculate the average acceleration ratios across all configurations for each dataset. When comparing on simpler dialogue datasets (i.e., ShareGPT and LMsys-1k), QSpec exhibits negligible throughput variation on multi-step reasoning tasks, particularly on coding tasks, despite a more pronounced performance decline of W4A4. Even on GSM8K and MATH tasks, a higher throughput is observed, due to the application of 8-shot and 4-shot prompts, respectively. This finding supports our observations in Sec. 2, where we attribute the reduction in multi-step reasoning capabilities to a few token changes that cause a worsening snowball effect, rather than numerous token prediction errors.
Larger models tend to yield better speedup ratios. We compare the average acceleration on all datasets across different models, and find a gradually-increasing acceleration as the base model scales up [^5] It is noteworthy that Llama2-7B shows higher speedup than Llama3-8B. This stems from the size difference primarily related to vocabulary, coupled with the introduction of Group-Query Attention (Ainslie et al., 2023), reducing the non-attention computation workload.. This overall upward trend indicates a promising outlook for our approach with larger models, although further experiments are needed for confirmation. Due to resource limitations, this will be addressed in future work.
Figure 4: Per-valid-token latency comparison of QSpec and W4A16 across different models and batch sizes. The latency of QSpec is further decomposed into draft and verify categories.
Latency Composition. As illustrated in Figure 4, we compute the per-valid-token latency by dividing the total latency by only the number of accepted tokens before averaged on all evaluation datasets. Notably, QSpec achieves remarkable latency savings ranging from 35.2% to 50.2%. Besides, the per-token latency is further decomposed into two components: draft and verify latency. Clearly, the primary gains of QSpec arise from the rapid drafting capability and the reduced latency achieved through the parallel verification of multiple tokens.
(a) Llama3.2-3b
(b) Llama3-8b-instruct
Ablation on Draft Token Length.  To assess parameter sensitivity, we vary the draft token lengths $\gamma$, the sole hyper-parameter of QSpec, from 2 to 7 across all the benchmarks using Llama3-8b-instruct and Llama3.2-3b models. As depicted in Figure 5, an increase in $\gamma$ leads to a gradual decline in the token acceptance rate, since all subsequent tokens are discarded once a token is rejected. Nevertheless, even at $\gamma=7$, the token acceptance rate remains relatively high, approximately $70\%$, compared to $28\sim 58\%$ in 160m-7b draft-target model pair under $\gamma=5$ in conventional speculative decoding (Liu et al., 2024). Additionally, a consistent improvement in throughput is observed compared to W4A16, indicating the robustness of QSpec with respect to $\gamma$.
5 Related work
Quantization is a common technique for deploying LLMs on resource-limited scenarios. Broadly, recent quantization algorithms can be classified into two categories: weight-only W4A16 and weight-activation joint W4A4. Notably, AWQ (W4A16) (Lin et al., 2024a) redistributes the quantization burden by scaling salient weight channels to protect them from degradation. In contrast, W4A4 aggressively quantizes activations to leverage low-precision hardware for improved speed at the cost of model quality degradation. To address this challenge, Atom (Zhao et al., 2024c) proposes reordering outlier channels in the activation through offline profiling. Similarly, QuaRot (Ashkboos et al., 2024) employs Hadamard matrices to apply computational invariance on weights. Despite these advancements, our observations indicate that W4A4 methods still exhibit substantial degradation compared to weight-only quantization approaches across multi-step reasoning tasks. On the other hand, adaptive quantization aims to optimize the trade-off between quantization-induced quality degradation and computational acceleration by mixed precision. LLM-PQ (Zhao et al., 2024a) proposes an adaptive layer-wise bitwidth selection approach, while QAQ (Dong et al., 2024) focuses on KV-cache bitwidth optimization. Other works operate at finer granularity to address outliers (Lee et al., 2024). However, these methods cannot fully recover the generation quality of higher precision.
Speculative Decoding leverages a draft model to generate candidate tokens, which are then validated by a target model (Leviathan et al., 2023). Recent research has primarily focused on improving the acceptance rate and generation speed of candidate tokens. SpecInfer (Miao et al., 2024) introduces a boost-tuned small language model to generate candidate tokens in tree structures, enabling single-pass verification. In contrast, EAGLE (Li et al., 2024) adopts an aggressive pruning strategy for the draft model’s architecture, allowing penultimate layer feature prediction with minimal computational overhead. Self-speculative decoding, a subset of this technique, employs a single model for both draft generation and verification. LayerSkip (Elhoushi et al., 2024) introduces a training methodology for early exit with layer drop, subsequently verifying partially generated tokens through full model inference. Medusa (Cai et al., 2024) constructs a generation tree of multiple candidate continuations by augmenting the original LLM with additional heads atop the final hidden state while relaxing the acceptance policy. However, these approaches inevitably require retraining of the original model, which can be computationally expensive and time-consuming.
Parameter Sharing
has been extensively applied for various purposes in previous research. Targeting parameter savings, Universal Transformer (Dehghani et al., 2018) shares all layers within a transformer model, while Subformer (Reid et al., 2021) shares its middle layers without sacrificing performance. Similarly, DictFormer (Lou et al., 2021) reparameterizes the model using a shared dictionary alongside unshared coefficients and indices, achieving reduced parameter redundancy and faster computations. Pires et al. (2023) enhances both accuracy and latency by implementing a single, larger shared feed-forward network across the encoder. In a different domain, Wang et al. (2024b; a) and Kopiczko et al. (2023) leverage parameter sharing in low-rank adaptation (LoRA) (Hu et al., 2021) to improve parameter efficiency. Unlike these methods, our focus is on sharing low-precision weights from two quantization schemes to maintain memory overhead.
6 Conclusion
In this paper, we begin by validating that multi-step reasoning tasks can capture performance degradation incurred by activation quantization more sensitively and consistently than current evaluation protocols, advocating for their incorporation for a more comprehensive assessment. With nearly cost-free execution switching and high token-level similarities, we introduce QSpec, a novel quantization paradigm that seamlessly synergizes two complementary weight-shared quantization schemes with speculative decoding. Empirically, QSpec achieve up to 1.80×1.80$\times$1.80 × acceleration without any quality sacrifice across diverse settings. Alongside consistent memory consumption and a plug-and-play property, these advantages distinguish QSpec from any existing solution, promising it for high-fidelity quantization deployment, particularly in memory-constrained scenarios.
Appendix A Evaluation Datasets
We utilize the following datasets for evaluation:
A.1 Language Modeling and Commonsense Reasoning
@@ A.4 Chatbot Datasets
  • LMsys-chat-1M-1K Zheng et al. (2023): Gathering one million authentic conversations with 25 leading large language models (LLMs), this dataset was sourced from over 210,000 unique IP addresses interacting with the Vicuna demo and Chatbot Arena websites.
Appendix B Datasets Sampling
@@
Appendix C Recommended Applications
We further analyze the best practice for efficiently deploying QSpec based on our empirical results:
Higher Matrix Multiplication Cost.
The primary advantage of low-precision quantization over weight-only quantization lies in their ability to better utilize non-attention compute kernels for speedup. However, certain architectural choices, such as Group-Query Attention (Ainslie et al., 2023) reduce the general matrix multiplication (GEMM) cost in the QKV (query, key, value) computations, which in turn diminishes the overall speedup of Llama3-8b compared to Llama2-7b.
Appropriate Batch Size.
The optimal performance of QSpec requires an appropriately sized batch. Specifically, the batch size should be neither too large nor too small. Excessively large batch sizes can cause the performance W4A16 verify to degrade as it scales proportionally with $\gamma$. Conversely, very small batch sizes are also suboptimal because low-precision kernels are most effective when computations are compute-bound; only memory footprint reductions offer benefits if computations are memory-bound. Moreover, excessive small batch size may not meet the tensor shape requirements necessary to exploit tensor cores (Zhao et al., 2024b) fully.
Additionally, in our experiments, we observed that some requests fail to generate the correct termination token and instead continue generating until they are truncated at the maximum generation length. As batch size increases, these excessively long requests are redundantly computed up to their maximum sequence length within the computing batch due to the first-in, first-out mechanism. This negative impact is magnified with larger batch sizes and draft-verify mechanism, reducing overall pipeline efficiency. This effect is particularly pronounced for the MBPP and HumanEval datasets, where the ratio of maximum length to average generation length is relatively large. Consequently, pipeline efficiency decreases further, leading to significant performance degradation with larger batches. Implementing appropriate early exit criteria would address this issue.
These observations suggest that QSpec is particularly well-suited for applications involving, non-attention compute-intensive and moderate batch sizes, where its advantages can be fully leveraged for optimal performance.

### v2

@@ A.1   Acceleration Evaluation
3B11footnotemark: 1 W16A16 8 511.1 588.7 756.6 647.2 785.7 711.2
16 666.5 845.6 1171.0 948.3 1292.2 1126.4
32 833.4 1081.5 1697.7 $1$.6 1975.6 1553.3
W4A4 8 804.7 921.2 1002.0 892.6 1091.6 990.3
16 1109.1 1374.5 1548.0 1289.8 1763.5 1581.0
32 1424.3 1899.3 2300.6 1488.2 2777.3 2194.4
W4A16 8 420.0 476.7 604.5 535.7 610.4 559.8
16 578.5 715.9 989.7 804.4 1080.2 925.8
32 726.3 933.8 1536.7 954.4 1704.5 1336.4
QSpec 8 594.1 (1.41$\times$) 648.2 (1.36$\times$) 760.1 (1.26$\times$) 723.6 (1.35$\times$) 787.5 (1.29$\times$) 738.8 (1.32$\times$) 1.33$\times$
16 811.5 (1.40$\times$) 936.0 (1.31$\times$) 1157.8 (1.17$\times$) 1042.1 (1.30$\times$) 1294.5 (1.20$\times$) 1171.4 (1.27$\times$) 1.27$\times$
32 1030.4 (1.42$\times$) 1240.2 (1.33$\times$) 1617.4 (1.05$\times$) 1248.5 (1.31$\times$) 1969.6 (1.16$\times$) 1576.0 (1.18$\times$) 1.24$\times$
Avg. 1.41$\times$ 1.33$\times$ 1.16$\times$ 1.32$\times$ 1.21$\times$ 1.25 $\times$ 1.28$\times$
7B W16A16 8 213.4 254.3 278.8 316.7 322.4 285.3
16 290.3 362.1 447.7 505.1 541.3 441.6
32 340.9 441.6 585.3 663.6 735.3 564.2
W4A4 8 349.5 411.7 396.1 471.2 471.8 419.4
16 496.6 612.2 614.3 749.5 760.9 642.6
32 620.0 793.6 801.5 1043.9 1083.2 865.5
W4A16 8 165.0 193.1 224.5 240.2 243.5 220.2
16 231.8 286.5 384.4 407.3 435.9 358.0
32 268.9 359.9 480.0 555.9 620.2 470.1
QSpec 8 253.7 (1.54$\times$) 291.5 (1.51$\times$) 298.3 (1.33$\times$) 350.9 (1.46$\times$) 345.7 (1.42$\times$) 310.3 (1.41$\times$) 1.44$\times$
16 359.8 (1.55$\times$) 420.2 (1.47$\times$) 466.7 (1.21$\times$) 555.2 (1.36$\times$) 557.8 (1.28$\times$) 473.1 (1.32$\times$) 1.37$\times$
32 441.8 (1.64$\times$) 527.2 (1.46$\times$) 575.3 (1.20$\times$) 749.4 (1.35$\times$) 770.0 (1.24$\times$) 628.4 (1.34$\times$) 1.39$\times$
Avg. 1.58$\times$ 1.48$\times$ 1.25$\times$ 1.39$\times$ 1.31$\times$ 1.36$\times$ 1.39$\times$
8B W16A16 8 189.4 211.5 256.0 259.1 290.7 265.8
16 262.0 311.2 408.7 401.2 511.0 447.4
32 303.8 390.8 566.3 522.6 820.0 649.8
W4A4 8 295.3 323.5 344.6 354.4 395.9 366.8
16 431.4 503.3 536.8 566.4 697.5 621.1
32 532.8 688.5 755.7 763.7 1167.9 956.8
W4A16 8 155.6 173.8 215.0 208.7 231.1 215.6
16 222.9 263.0 354.8 345.9 422.8 369.4
32 $\dagger$ $\dagger$ 509.8 468.7 706.0 580.5
QSpec 8 222.6 (1.43$\times$) 233.9 (1.35$\times$) 256.7 (1.19$\times$) 271.5 (1.30$\times$) 285.0 (1.23$\times$) 268.3 (1.24$\times$) 1.29$\times$
16 322.6 (1.45$\times$) 362.5 (1.38$\times$) 402.7 (1.14$\times$) 438.5 (1.27$\times$) 507.5 (1.20$\times$) 453.5 (1.23$\times$) 1.28$\times$
32 400.2 ($\dagger$) 362.5 ($\dagger$) 578.1 (1.13$\times$) 573.0 (1.22$\times$) 798.8 (1.13$\times$) 684.5 (1.18$\times$) 1.27$\times$
Avg. 1.44$\times$ 1.36$\times$ 1.15$\times$ 1.26$\times$ 1.19$\times$ 1.22$\times$ 1.27 $\times$
13B11footnotemark: 1 W16A16 8 121.9 146.6 183.1 182.0 187.1 160.1
16 169.6 211.2 304.4 291.0 311.0 243.0
32 202.4 253.8 426.0 423.5 311.0 334.2
W4A4 8 194.7 228.2 253.6 261.5 259.8 228.2
16 288.3 349.2 415.3 424.9 431.5 348.4
32 369.8 469.9 606.7 665.4 431.5 508.8
W4A16 8 94.8 112.9 143.4 140.0 146.7 127.9
16 136.1 171.9 250.8 236.9 255.9 207.2
32 $\dagger$ $\dagger$ 376.4 365.5 255.9 287.4
QSpec 8 148.2 (1.56$\times$) 167.9 (1.49$\times$) 193.6 (1.35$\times$) 201.2 (1.44$\times$) 194.5 (1.33$\times$) 174.0 (1.36$\times$) 1.42$\times$
16 212.8 (1.56$\times$) 248.6 (1.45$\times$) 316.8 (1.26$\times$) 323.3 (1.36$\times$) 327.4 (1.28$\times$) 266.9 (1.29$\times$) 1.29$\times$
32 266.6 ($\dagger$) 320.0 ($\dagger$) 451.5 (1.20$\times$) 483.0 (1.32$\times$) 327.4 (1.28$\times$) 379.3 (1.32$\times$) 1.32$\times$
Avg. 1.56$\times$ 1.47$\times$ 1.27$\times$ 1.37$\times$ 1.29$\times$ 1.32$\times$ 1.38$\times$
Table 7 presents the comprehensive comparison of token generation throughput across multiple dimensions: model sizes, quantization configurations, and batch sizes, evaluated on various datasets. It is noteworthy that Llama2-7B shows higher speedup than Llama3-8B. This stems from the size difference primarily related to vocabulary, coupled with the introduction of Group-Query Attention (Ainslie et al., 2023), reducing the computation workload.
Versus speculative decoding. Table 8 presents a comprehensive comparison between our approach and the EAGLE method across multiple benchmarks.
Table 8: Performance comparison of EAGLE-Quant, QSpec, W4A16, and W4A4 on Llama-2-7b-chat-hf across different batch sizes. Results are reported for GSM8K (8-shot), MATH (4-shot), MBPP (0-shot), HumanEval (0-shot), ShareGPT, and LMsys-1k benchmarks. “OOM” indicates out-of-memory errors. Better case for QSpec or EAGLE is marked in gray. In the case of batch size=8, the speedup ratio of QSpec compared to EAGLE is indicated in parentheses next to the data points.
Method Batch Size GSM8K (8-shot) MATH (4-shot) MBPP (0-shot) HumanEval (0-shot) ShareGPT LMsys-1k
EAGLE 1 65.81 70.11 68.53 49.15 79.60 71.29
8 140.16 210.52 138.01 136.86 247.42 167.57
16 OOM OOM OOM OOM OOM OOM
1 51.25 48.36 56.87 54.22 56.77 56.14
QSpec 8 208.95 (1.49$\times$) 249.97 (1.19$\times$) 185.13 (1.34$\times$) 185.99 (1.36$\times$) 329.44 (1.33$\times$) 260.48 (1.55$\times$)
16 292.82 356.93 269.87 255.11 562.07 463.35
W4A16 1 59.80 63.65 75.49 72.04 76.04 72.27
8 146.34 185.52 180.12 163.54 250.06 213.66
16 190.09 251.59 254.24 211.49 458.57 371.38
W4A4 1 64.79 66.32 74.72 73.47 73.09 71.55
8 284.84 369.27 250.11 256.54 492.21 393.12
16 401.77 540.82 357.62 330.71 895.77 713.67
A.2   Ablation Studies
Comprehensive Performance Trade-offs: QSpec versus W4A16
We conduct a thorough analysis of the trade-offs between throughput and accuracy for our proposed framework against all baseline implementations. Figure 7 illustrates this comparison, plotting generation quality (accuracy) against computational efficiency (throughput). Our analysis reveals that while W4A4 suffers substantial performance degradation (18.5-39.5% reduction) on multi-step reasoning benchmarks compared to W4A16, QSpec achieves comparable accuracy to W4A16 while delivering significantly higher throughput. Although QSpec’s accuracy is marginally lower than W16A16 due to weight quantization-induced memory optimization, it successfully preserves the performance characteristics of W4A16 while offering superior computational efficiency.
Refer to caption Batch Size 8     Refer to caption Batch Size 16
Figure 7: Comparison of accuracy and efficiency among W16A16, W4A16, W4A4, and QSpec across various datasets with batch sizes of 8 and 16 on Llama3-8b-instruct model. The bars and lines represent the accuracy and throughput of each method.
A.3   Datasets and Sampling
For the PIQA and Winogrande datasets, we randomly select 500 questions from each for performance evaluation. In contrast, we process the entire GSM8K and MATH datasets, as detailed in Atom Zhao et al. (2024b). When adapting QSpec to the Quarot method, we sample 200 samples from the GSM8K dataset, while sampling 700 from the MATH dataset, ensuring balanced representation by selecting 100 questions from each of the seven distinct question types. Additionally, we sample 200 questions from the MBPP dataset, while processing the entire HumanEval dataset. This preserves a thorough assessment of the model’s performance across various datasets.
For acceleration evaluation, we maintain the random seed at 42, and sample 100 samples from the GSM8K, MATH, MBPP, HumanEval, ShareGPT, and LMSys datasets, respsectively. This consistent methodology guarantees that our evaluation remains reproducible and representative across these diverse datasets.
Appendix B Evaluation Datasets
We comprehensively evaluate the performance of QSpec, covering language modeling, commonsense reasoning, mathematical reasoning, code and chatbot. Each dataset is briefly introduced below, along with the associated sampling strategy for enhanced efficiency.
Language Modeling and Commonsense Reasoning.
@@ Appendix B Evaluation Datasets
  • LMsys-chat-1M-1K Zheng et al. (2023): Gathering one million authentic conversations with 25 leading large language models, this dataset was sourced from over 210,000 unique IP addresses interacting with the Vicuna demo and Chatbot Arena websites.
Appendix C Datasets Sampling